OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (2, 5) and (–1, 8)? y + 5 = 1(x + 2) y + 8 = –1(x – 1) y – 5 = –1(x – 2) y – 8 = 1(x – 2) What is the equation of the line in standard form that passes through the point (3, –2) and is parallel to the line y = –4x + 1? x + 4y = 10 x – 4y = –10 4x – y = –10 4x + y = 10

OpenStudy (helder_edwin):

do u know how to compute the slope between two points??

OpenStudy (anonymous):

not very well no. could you please explain?

OpenStudy (helder_edwin):

the formula is \[ \large \frac{y_0-y_1}{x_0-x_1} \] for given points \((x_0,y_0)\) and \((x_1,y_1)\)

OpenStudy (helder_edwin):

u have (2,5) and (-1,8) try it!!

OpenStudy (anonymous):

i'm completely stuck. could you walk through the problems for me?

OpenStudy (helder_edwin):

OK. the slope would be \[ \large m=\frac{8-5}{-1-2}=\frac{3}{-3}=-1 \] got it???

OpenStudy (anonymous):

Okay, i got that part.

OpenStudy (helder_edwin):

OK. this is the slope of the line u r looking for.

OpenStudy (helder_edwin):

the line in slope-point form is \[ \large y-y_0=m(x-x_0) \]

OpenStudy (anonymous):

Right, i know that part.

OpenStudy (helder_edwin):

u already have m=-1 and two points (-1,8) and (2,5). u can choose either one of the two points!!! choose one!!

OpenStudy (anonymous):

Okay, (-1,-8)

OpenStudy (helder_edwin):

no (-1,8)

OpenStudy (helder_edwin):

then u have \(m=-1\) and the point \((-1,8)=(x_0,y_0)\). replace this into \[ \large y-y_0=m(x-x_0) \]

OpenStudy (anonymous):

okay, did that.

OpenStudy (helder_edwin):

show me!!

OpenStudy (anonymous):

I'm doing it on my homework. But, i think i just figured out what i did for the first one. what did you get?

OpenStudy (helder_edwin):

ok it should be \[ \large y-8=-1(x-(-1)) \] or \[ \large y-8=-(x+1) \]

OpenStudy (anonymous):

thank you so much for your help. what did you get for the second one?

OpenStudy (helder_edwin):

u have the line \[ \large y=-4x+1 \] it has the form slope-intercept \[ \large y=mx+b \] so m=-4

OpenStudy (anonymous):

okay

OpenStudy (helder_edwin):

it just as above. u have m=-4 and one point (3,-2). do what u did before but this time do the algebra.

OpenStudy (anonymous):

Yeah thats where i lack. Okay, thank you for your help.

OpenStudy (helder_edwin):

u r welcome